# CaTS-Bench: Can Language Models Describe Numeric Time Series?

### v1

@@ 4.2 Q&A Tasks
No model consistently dominates across all categories. Models handle binary-choice time series comparisons better, likely due to the narrower range of options. Matching a time series to a caption is harder than the reverse, and plot matching is the most challenging, highlighting a key VLM weakness: linking numeric patterns with visual features. Proprietary models (GPT-4o, Gemini 2.0 Flash) lead, while among open-source models, Phi-4 M.I. excels in time series and statistical reasoning. Finetuning on TSC yields mixed results: some models (e.g., Phi-4 M.I., Idefics 2) gain in specific sub-tasks, while others drop in performance. Notably, finetuning often fails to improve Q&A accuracy, likely due to task misalignment and catastrophic forgetting. As Table 17 shows, humans achieve the highest overall scores, though top models sometimes outperform them on distraction-prone tasks. Notably, all models perform near-random on plot matching, whereas humans score nearly perfectly. Despite the tasks’ apparent simplicity, they reveal fundamental limitations in VLMs’ temporal reasoning capabilities which suggests the need to address basic time series understanding before tackling more complex applications.
@@
(b) Pretrained VLMs
(c) Finetuned VLMs
Figure 3: Model accuracy across Q&A sub-tasks. Proprietary models perform best, pretrained models lag behind, and finetuned models struggle across all tasks.
@@ 4.3 Role of the Visual Modality
We perform a modality removal experiment by stripping away the time series plot and providing only the associated textual metadata and the numeric values of the time series. This quantifies the contribution of the visual channel and enables a better understanding of the model’s captioning performance. We evaluate a selected subset of pretrained baselines to assess their intrinsic reliance on vision. Full results can be found in Appendix I.1.
Our experiments suggest that the additional contribution of the visual modality to caption quality is insignificant for most models. As shown in Figure 4, most models show only marginal performance drops, or even slight gains, when the time series plot is removed, suggesting a strong dependence on textual priors over visual understanding. In particular, models such as Idefics2, Phi-4 M.I., and InternVL perform better in text-only settings on most metrics, hinting that generation is largely driven by language pretraining or instruction tuning rather than true visual interpretation.
Figure 4: Performance deltas between VL (vision-language input) and L (text-only input). Each cell shows $\Delta=\text{VL}-\text{L}$. Blue indicates better performance due to the visual input; red the opposite.
Models such as QwenVL, LLaVA 1.6 and Claude 3 Haiku maintain strong performance with visual input, but the performance gap ($\Delta$) remains modest, underscoring the underuse of plot-based information. Interestingly, the numeric score tends to decline when visual input is removed, hinting at weak but present reliance on the plot for numeric reasoning. These results point to a subtle yet important misalignment: models are exposed to visual data but often fail to meaningfully reason with it. This phenomenon is not limited to line plots, as discussed in I.3, even more expressive visual forms (e.g., Gramian Angular Fields and recurrence plots) fail to trigger visual reasoning of current VLMs in TSC.
@@
To better understand how VLMs process plots during caption generation, we examined their attention maps. The analysis revealed minimal visual grounding: models concentrated predominantly on textual elements in the plots (e.g., axis labels and titles), with limited evidence of attending to the actual line trends. Attention to visual patterns was sporadic, weak, and inconsistent, suggesting that learned parameters largely disregard visual cues in favor of textual priors. This qualitative evidence highlights the gap between nominal multimodal input and actual integration. Full results are reported in Appendix I.2 and Figure 7.
It is important to note that the under-utilization of visual inputs observed in our experiments is not a limitation of CaTS-Bench itself, but rather a reflection of current VLM capabilities. The benchmark explicitly provides both time series plots and rich metadata, creating ample opportunity for multimodal reasoning. That most models default to textual priors instead of leveraging visual signals highlights a critical gap in the field. We view this as an opportunity for future research: developing models that better integrate plot-based information with textual and numeric cues to advance the broader goal of genuine multimodal understanding in time series analysis.
@@ 5 Conclusion
We introduced CaTS-Bench, the first large-scale, multimodal benchmark for context-aware time series captioning and reasoning. Built from $11$ diverse real-world datasets, it combines numeric series, metadata, visual plots, and validated captions to provide a challenging testbed beyond synthetic or narrow benchmarks. A key contribution is not only the benchmark itself, but also the scalable data curation pipeline we developed to generate high-quality captions. This pipeline leverages an oracle LLM anchored in metadata, rigorous verification through factual checks, diversity analyses, and a complementary human-revisited subset, making it both scalable and extensible to new domains. Our evaluation of leading VLMs revealed both progress and limitations. Finetuning greatly improves open-source models, enhancing fluency and numeric fidelity, while proprietary models show stronger performance overall. A consistent weakness lies in multimodal grounding: models largely ignore visual inputs, with plot matching emerging as the most difficult task. These findings reveal a critical gap in multimodal alignment and point toward the urgent need for models that can genuinely integrate numeric, textual, and visual cues. By releasing CaTS-Bench together with its evaluation suite, we provide the community with not only a rigorous foundation for advancing time series reasoning, but also a practical methodology for generating reliable, context-rich captions at scale, paving the way for more robust multimodal understanding in the future.
Ethical Statement
@@
The development of CaTS-Bench was guided by a commitment to ethical research practices. All datasets used in this work are publicly available and do not contain personally identifiable information (PII). The domains, such as climate, public health, and agriculture, were chosen for their public relevance and data accessibility. Our use of an oracle LLM to generate semi-synthetic reference captions was a deliberate design choice to ensure scalability, particularly for a subjective task like captioning, where a single ground-truth is ill-defined. We have taken extensive measures to validate the quality, factual accuracy, and diversity of these semi-synthetic captions, as detailed in Section 3.2 and Appendix H, to mitigate the risk of propagating systemic biases from the oracle model. Our human-revisited test set is also an attempt to further ensure evaluation reliability. For our human evaluation studies, all participation was voluntary. We obtained informed consent from all participants, who were university students. The study’s purpose was clearly communicated, and all responses were collected anonymously to protect participant privacy, as shown in an example of a consent form in Appendix O.
Reproducibility Statement
To ensure the reproducibility of our research, all components of our work will be made publicly available upon publication.
  1. 1.
    Data: The complete CaTS-Bench dataset, including the numeric time series, metadata, generated plots, oracle-generated and human-revisited captions, and the diagnostic Q&A suite, are released at https://huggingface.co/datasets/a9f3c7e2/CaTSBench.
  2. 2.
    Code: We will release the source code for the entire data curation pipeline, model finetuning scripts, and the evaluation suite. The code will be hosted in a public repository to allow for complete replication of our results and to facilitate future research.
  3. 3.
    Models and Environment: All open-source models used in our experiments are explicitly named with version details provided in Appendix E. For proprietary models, we specify the exact model endpoints used at the time of the experiments. Detailed finetuning hyperparameters and hardware specifications are documented in Appendix D.
  4. 4.
    Evaluation: Our evaluation protocol relies on standard, well-established linguistic metrics and novel metrics that are precisely defined in F. All prompts used for caption generation, quality verification, and LLM-based scoring are provided in Appendix N to ensure that our evaluation can be replicated consistently. Furthermore, we conducted a robustness check (Appendix H.5), which demonstrated minimal variance across multiple runs, confirming the stability of our results.
@@
    LLMs were employed as data generators, producing semi-synthetic captions that serve as ground truth references in CaTS-Bench.
@@
Appendix A Limitations and Future Work
While CaTS-Bench represents a significant step toward multimodal, context-aware time series understanding, it also has limitations that suggest clear avenues for future work. First, the majority of captions are semi-synthetic and generated by a single oracle model (Gemini 2.0 Flash). Our validation studies confirm their factual reliability and linguistic diversity, with no clear evidence of bias; however, reliance on a single oracle may still introduce subtle, hidden modeling biases. Although we release a subset of data paraphrased by different LLMs, its limited scale points to future iterations of CaTS-Bench where semi-synthetic captions are fully generated from a broader pool of LLMs, thereby reducing such biases and better reflecting diversity in expression. Second, although we provide a curated human-revisited subset of captions, the scale of fully human-authored or expert-verified content remains limited. We acknowledge that captions written by human experts are often more insightful, and future work should incorporate such expert-written captions to enhance depth and interpretability. At the same time, human authors introduce their own stylistic biases, which should be considered when designing and evaluating the benchmark. Expanding this component, potentially by involving multiple domain experts in economics, healthcare, or climate data, would further strengthen the benchmark’s robustness and credibility.
Overall, we view CaTS-Bench as a scalable foundation rather than a fixed resource, with ample room to grow through multi-oracle captioning, richer human input, and extended coverage of temporal reasoning tasks.
@@ Appendix B Source Datasets
    Air Quality – Hourly air pollution data from 453 Indian cities (2010–2023), covering 30+ parameters including PM2.5, NOx, CO, and SO2, compiled from CPCB Jha (2023).
@@
    Border Crossing – Monthly inbound border crossing counts at U.S.-Mexico and U.S.-Canada ports, disaggregated by transport mode and collected by U.S. Customs and Border Protection U.S. Department of Transportation (n.d.).
@@
    Crime – Incident-level crime reports in Los Angeles from 2020 onward, provided by LAPD OpenData and updated biweekly, including NIBRS-compliant records of Los Angeles (n.d.).
@@
    Demography – Annual global indicators from the UN and World Bank (2000–2021) covering population growth, fertility, life expectancy, death rates, and median age to assess patterns of demographic change and collapse Aziz (1985).
@@
    Injury – Annual counts of fatal and severe road traffic injuries in California (2002–2010), disaggregated by transport mode and geography, from CDPH’s Healthy Communities Indicators of Public Health (n.d.).
@@
    COVID – Global daily COVID-19 case and death counts (2020), compiled by ECDC, covering over 200 countries with population-adjusted metrics European Centre for Disease Prevention and Control (2024).
@@
    CO2 – National-level per capita CO2 emissions and GDP trends from Our World in Data, adjusted for trade (consumption-based), spanning 1990–2023 Ritchie (2021).
@@
    Calories (Diet) – Food supply and caloric intake patterns from FAO Food Balance Sheets Food and Agriculture Organization of the United Nations (2024).
@@
    Walmart – Weekly sales data from 45 Walmart stores (2010–2012), enriched with features like temperature, fuel price, CPI, unemployment rate, and holiday flags Hassan (2020).
@@
    Retail – Transactional records from a UK-based online gift retailer (2010–2011), capturing item-level purchases, cancellations, and customer behavior Chen (2015).
@@
    Agriculture – Annual agricultural total factor productivity (TFP) indices from USDA for 1961–2022, covering outputs and inputs like land, labor, capital, and materials across countries USDA Economic Research Service (2024).
@@ Appendix D Hardware and Settings
Table 6: Configurations
@@
Grad. Acc. 12
Epochs 3
@@
Scheduler Cosine
@@
Dropout 0.05
Image res. 224–560
@@ Appendix E Baseline Models
We evaluate Gemini 2.0 Flash and Gemini 2.5 Pro Preview (Team et al., 2023), Claude 3 Haiku and Claude 3.7 Sonnet (Anthropic, 2024), GPT-4o (Achiam et al., 2023), InternVL 2.5 (8b & 38b) (Chen et al., 2024b), LLaVA v1.6 Mistral 7b (default) and 34b (Liu et al., 2023), Phi-4 Multimodal Instruct 5.6b (Abdin et al., 2024), Idefics 2 (8b) (Laurençon et al., 2024), SmolVLM (2b) (Marafioti et al., 2025), QwenVL (7b) (Bai et al., 2023), Llama 3.2 Vision (11b) (Grattafiori et al., 2024), and Gemma 3 (12b & 27b) (Team et al., 2025) for both TSC and Q&A tasks.
TSC requires precise numeric reasoning alongside text generation, making it suitable for program-aided language (PAL) models (Gao et al., 2023). Hence, we also evaluate QwenVL 32b by prompting it to generate a Python program that outputs the full time series caption. The program is executed in Python, and its return value is taken as the caption. Most ( 90%) generated programs succeed on the first attempt; if a program fails, we increase the token limit and regenerate until successful. The full prompt example can be found in Appendix N.4.

### v2

@@ 4.2 Results on Q&A Tasks
Models show no consistent winner across tasks. Binary time-series comparisons are easier, likely due to the narrower decision space, while caption matching is harder, and plot matching is hardest. Proprietary models lead, and Phi-4 M.I. is the strongest open-source model overall. Humans achieve the highest overall scores and are nearly perfect on plot matching, where all models remain near random. Despite the tasks’ apparent simplicity, they reveal fundamental limitations in VLMs’ temporal reasoning capabilities which suggests the need to address basic time series understanding before tackling more complex applications.
@@
(b) Open-source VLMs
Figure 3: Model accuracy across Q&A sub-tasks (except Temporal Matching). Proprietary models perform best, open-source models lag behind in all tasks.
Figure 4: MCQ accuracy (%) for exact temporal value lookup at different query positions. Short_Mid2 denotes sequences of length 6; other columns use long time series with queries at different temporal offsets. Results are averaged over 75 questions per temporal bucket.
Temporal matching also unveils significant weaknesses, with results shown in Fig. 4. Models perform near chance on long sequences ($\approx 20$$35\%$), with accuracy degrading as the queried time moves further from the series start in most models, while short sequences (length =6=6) yield higher accuracy ($\approx 50$$57\%$). These results indicate that current language models struggle with exact index-based temporal reasoning from text, and increased model scale does not reliably mitigate this limitation. For Short_Mid2 sequences, vision modestly improves accuracy over text-only input (Table 20), but precise temporal indexing remains challenging. See Appendix I.1 for more information about temporal buckets, baseline, and detailed results.
@@ 4.3 Role of the Visual Modality
We perform a modality removal experiment by stripping away the time series plot and providing only the associated textual metadata and the numeric values of the time series. This quantifies the contribution of the visual channel and enables a better understanding of the model’s captioning performance. We evaluate a selected subset of pretrained baselines to assess their intrinsic reliance on vision.
Our experiments show that visual input contributes little to caption quality for most models. As Figure 5 illustrates, removing the plot image produces only minor changes in both semantic and numeric performance, with numeric accuracy degrading modestly and sometimes even improving without the plot. This suggests that models rely mainly on textual priors rather than grounded interpretation of the plotted signal, so caption generation is driven primarily by language-level expectations about typical time-series behaviors.
The weak sensitivity to image removal indicates that visual input is only weakly used during generation, for both semantic and numeric metrics, implying that vision plays a limited role in numeric grounding. As discussed in Section H.2, this phenomenon is not restricted to line plots: more expressive encodings such as Gramian Angular Fields and recurrence plots likewise fail to elicit strong visual reasoning from current VLMs in TSC.
Figure 5: Performance change when adding visual input. Blue indicates improvement; red indicates degradation.
Figure 6 isolates the effect of numeric annotations on the plot, while keeping the visual modality fixed. In contrast to the full-versus-no-plot comparison, the impact here is substantially larger. When time-series values are explicitly annotated, models show clear improvements in fine-grained numeric metrics, particularly Max, Min, and Mean. Removing these annotations while retaining the plotted curve leads to pronounced drops in numeric fidelity. This pattern indicates that models benefit primarily from explicit numeric text embedded in the image, rather than from interpreting the visual shape of the time-series curve. The relatively small and inconsistent changes in Std further suggest that dispersion or variability is not reliably inferred from visual structure alone.
Figure 6: Performance deltas on statistical inference with and without annotations. Blue indicates improvement; red degradation.
@@
To better understand how VLMs use plots during caption generation, we examine their attention maps and find minimal visual grounding: models focus mainly on textual elements (e.g., axis labels and titles) with limited attention to the actual line trends. Attention to visual patterns is weak and inconsistent, suggesting that learned parameters largely ignore visual cues in favor of textual priors, highlighting a gap between nominal multimodal input and actual integration (full results in H.1 and Figure 9).
These findings indicate that current VLM pretraining paradigms are ill-suited for structured visual reasoning over time series plots: despite strong performance on natural images, their objectives align poorly with abstract, low-level visual encodings that require relational, temporal, and quantitative interpretation rather than object recognition. Consequently, models default to textual priors and explicit numeric cues, largely ignoring visual structure, even though CaTS-Bench explicitly provides both plots and rich metadata to support multimodal reasoning. That most models fail to leverage visual signals exposes a critical gap in current VLM capabilities and motivates future work on architectures that more effectively integrate plot-based information with textual and numeric cues.
@@ 5 Conclusion
We presented CaTS-Bench, a multimodal benchmark for TSC and time-series reasoning that emphasizes linguistic expressiveness, numeric fidelity, and grounded multimodal alignment. By combining a human-rewritten gold standard with a scalable synthetic augmentation pipeline, CaTS-Bench enables rigorous evaluation of free-form narrative generation over time series while addressing data scarcity. Our results show that proprietary VLMs provide strong zero-shot baselines, while finetuning open-source models on synthetic data yields substantial gains in TSC. However, zero-shot VLMs exhibit systematic failures such as hallucinated statistics in free-form captioning. Diagnostic Q&A tasks further reveal persistent weaknesses in exact temporal indexing and multimodal integration, indicating that the current VLM pretraining paradigm is ill-designed for time series reasoning. Overall, CaTS-Bench exposes key limitations in grounded time series understanding and provides a foundation for developing more reliable multimodal generative models in numeric domains.
Limitations
While CaTS-Bench represents a significant step toward multimodal, context-aware time series understanding, it also has limitations that suggest clear avenues for future work.
First, the training captions are synthetic and generated by a single oracle model (Gemini 2.0 Flash). Our extensive validation studies confirm their factual reliability and linguistic diversity; however, reliance on a single oracle may still introduce subtle modeling biases during finetuning. As discussed earlier, this design choice is intended as a proof of concept to demonstrate the feasibility and effectiveness of scalable augmentation. Extending the pipeline to multiple oracles is straightforward but beyond the scope of this work. Although we release a small subset of paraphrases generated by different LLMs, its limited scale highlights opportunities for future extensions.
Second, although we provide a human-rewritten test set, the scale of expert-verified content remains limited. We acknowledge that captions written by domain experts can offer deeper insights, and future work could incorporate expert-authored annotations across domains such as economics, healthcare, or climate data. At the same time, human authors introduce their own stylistic biases, which should be considered when designing and evaluating benchmarks.
Overall, we view CaTS-Bench as a scalable foundation rather than a fixed resource, with ample room to grow through richer human input, multi-oracle augmentation, and extended coverage of temporal reasoning tasks.
Ethical Considerations
@@
The development of CaTS-Bench was guided by a commitment to ethical research practices. All datasets used in this work are publicly available and do not contain personally identifiable information (PII). The domains, such as climate, public health, and agriculture, were chosen for their public relevance and data accessibility. Our use of an oracle LLM for data augmentation was a deliberate design choice to ensure scalability. We have taken extensive measures to validate the quality of the synthetic captions, gauging their factual accuracy and diversity, as detailed in Section 3.3 and Appendix G. Our human-rewritten test set is also an attempt to further ensure evaluation reliability. For our human Q&A evaluation studies, all participation was voluntary. We obtained informed consent from all participants, who were university students. The study’s purpose was clearly communicated, and all responses were collected anonymously to protect participant privacy, as shown in an example of a consent form in Appendix M.
@@
    LLMs were employed as data generators, producing synthetic captions that mainly serve as training data.
@@
Appendix A Motivation and Utilities of TSC
An interesting question to think about is whether TSC is a meaningful task at all, given that many properties of a time series (e.g., means, extrema, trends) can be computed exactly using deterministic algorithms or statistical tools. While such tools are indeed precise, they fundamentally address a different problem. TSC is not concerned with computing statistics, but with generating free-form natural language descriptions that summarize, contextualize, and communicate temporal data in a human-interpretable way.
In many real-world settings, users do not interact directly with raw statistics or code outputs, but instead rely on narrative descriptions. Business reports, financial dashboards, scientific summaries, ECG reports, policy briefs, and automated analytics systems routinely present trends and patterns through textual explanations accompanying plots. In these contexts, the goal is not merely to compute correct values, but to decide what is salient, how to describe it, and how to express uncertainty, comparison, or change over time in natural language. These decisions are inherently under-specified and context-dependent, making them unsuitable for purely deterministic pipelines.
LLMs are therefore a natural candidate for this task, as they are designed to produce fluent, context-aware descriptions. However, our results demonstrate that current models struggle precisely at this interface between numeric data and free-form generation. We observe frequent hallucination of numeric values, incorrect temporal alignment between values and timestamps, and failures in simple comparative reasoning. Moreover, when strong textual cues are present, models often ignore the underlying plot, generating plausible but weakly grounded descriptions driven by language priors rather than data. These failure modes are not exposed by deterministic statistical tools, which are guaranteed to be correct by construction, but are critical in generative settings where outputs are unconstrained and must be trusted by downstream users. The fact that calculators or statistical algorithms can compute accurate summaries does not obviate the need to evaluate how models describe those summaries in natural language. Instead, it highlights the importance of studying TSC as a distinct problem, one that probes a model’s ability to integrate numeric reasoning, temporal structure, and linguistic generation.
From this perspective, TSC serves as a diagnostic task that reveals the limits of current LLMs and VLMs in grounded free-form generation. Our findings do not argue against deterministic methods, but rather show that replacing human-written narrative descriptions with model-generated captions with direct prompting remains challenging. Understanding these limitations is essential for developing reliable systems that can safely and effectively communicate temporal data in real-world settings.
@@ Appendix B Source Datasets
    Air Quality – Hourly air pollution data from 453 Indian cities (2010–2023), covering 30+ parameters including PM2.5, NOx, CO, and SO2, compiled from CPCB jha2023airquality.
@@
    Border Crossing – Monthly inbound border crossing counts at U.S.-Mexico and U.S.-Canada ports, disaggregated by transport mode and collected by U.S. Customs and Border Protection bts_bordercrossing.
@@
    Crime – Incident-level crime reports in Los Angeles from 2020 onward, provided by LAPD OpenData and updated biweekly, including NIBRS-compliant records lapd_crime_2020.
@@
    Demography – Annual global indicators from the UN and World Bank (2000–2021) covering population growth, fertility, life expectancy, death rates, and median age to assess patterns of demographic change and collapse aziz1985population.
@@
    Injury – Annual counts of fatal and severe road traffic injuries in California (2002–2010), disaggregated by transport mode and geography, from CDPH’s Healthy Communities Indicators cdph_roadtrafficinjuries.
@@
    COVID – Global daily COVID-19 case and death counts (2020), compiled by ECDC, covering over 200 countries with population-adjusted metrics ecdc_covid_data.
@@
    CO2 – National-level per capita CO2 emissions and GDP trends from Our World in Data, adjusted for trade (consumption-based), spanning 1990–2023 owid-co2-gdp-decoupling.
@@
    Calories (Diet) – Food supply and caloric intake patterns from FAO Food Balance Sheets fao_diet_fbs.
@@
    Walmart – Weekly sales data from 45 Walmart stores (2010–2012), enriched with features like temperature, fuel price, CPI, unemployment rate, and holiday flags kaggle_walmart.
@@
    Retail – Transactional records from a UK-based online gift retailer (2010–2011), capturing item-level purchases, cancellations, and customer behavior online_retail_352.
@@
    Agriculture – Annual agricultural total factor productivity (TFP) indices from USDA for 1961–2022, covering outputs and inputs like land, labor, capital, and materials across countries usda_agriculture_productivity.
@@ Appendix D Hardware and Settings
Table 5: Fine-tuning configuration details
@@
Gradient Accumulation Steps 12
Training Epochs 3
@@
Learning Rate Scheduler Cosine
@@
Dropout Probability 0.05
Input Image Resolution 224–560
@@ Appendix E Baseline Models
For TSC, we evaluate Gemini 2.0 Flash (team2023gemini), Claude 3 Haiku (TheC3), GPT-4o (achiam2023gpt), InternVL 2.5 (8b & 38b) (chen2024expanding), LLaVA v1.6 Mistral 7b (default) and 34b (liu2023llava), Phi-4 Multimodal Instruct 5.6b (abdin2024phi), Idefics 2 (8b) (laurenccon2024matters), SmolVLM (2b) (marafioti2025smolvlm), QwenVL (7b) (Qwen-VL), Llama 3.2 Vision (11b) (grattafiori2024llama), and Gemma 3 (12b & 27b) (team2025gemma) for both TSC and Q&A tasks.
TSC requires precise numeric reasoning alongside text generation, making it suitable for program-aided language (PAL) models (gao2023PAL). Hence, we also evaluate QwenVL 32b by prompting it to generate a Python program that outputs the full time series caption. The program is executed in Python, and its return value is taken as the caption. Most ( 90%) generated programs succeed on the first attempt; if a program fails, we increase the token limit and regenerate until successful. The full prompt example can be found in Appendix L.4.